OpenStudy (anonymous):

18. For the function rule y = -2x - 3, provide the x-y coordinates of the y-intercept and the x-intercept.

OpenStudy (imstuck):

The x coordinate is where y = 0. Replace y with 0 and solve for x. The y intercept coordinate is where x = 0, so replace x with 0 and solve for y. Very simple math.

OpenStudy (imstuck):

\[y = -2(0)-3\]\[y=-3\]\[0=-2x-3\]\[3=-2x\]\[-\frac{ 3 }{ 2 }=x\]